# Sharp Self-Normalized Concentration Inequalities of Marginal Mean with Sample Variance Only

### v2

@@ A.2 Proof of Results in Section 2
Step 2 (Generic high-probability bound) According to the proof of Theorems 2 and 3, we have generic high-probability (probability larger than $1-2\delta$) bound as follow,
$|S_{nn}(h_{n})|\leq|S_{nb_{m}}(h_{n})|+|r_{n}|=|\sum_{j=1}^{m}M_{nj}|+|r_{n}|$
$\leq$ $\max\left\{\sqrt{2\log(\frac{1}{\delta})}\left(\sqrt{\sum_{j=1}^{m}T^{2}_{nj}}+\sqrt{\sum_{j=1}^{m}R^{2}_{nj}}+\sqrt{\xi_{n}}\right),2\log(\frac{1}{\delta})(\max_{1\leq j\leq m}|T_{nj}|+\max_{1\leq j\leq m}|R_{nj}|)\right\}+|r_{n}|.$
Then, some simple algebra shows that, for any given $\delta>0$ and arbitrary $\xi_{n}\searrow 0$, we have
$\mathbb{P}\Big(|\frac{1}{n}S_{nn}(h_{n})|\leq\tilde{\nu}_{n}(\delta)\max\{\mathbf{A},\mathbf{B}\}+|\frac{r_{n}}{n}|\Big)\geq 1-2\delta,$ (A.38)
$\mathbf{A}=$ $\sqrt{\frac{2\log(\frac{1}{\delta})[\sum_{j=1}^{m}(\sum_{i\in B_{j}}h_{n}(Z_{ni})-\bar{H}_{n})^{2}]}{n^{2}}}+\frac{\max_{1\leq j\leq m}|R_{nj}|}{\sqrt{n[l]}}\sqrt{2\log(\frac{1}{\delta})}+\sqrt{\frac{2\xi_{n}\log(\frac{1}{\delta})}{n^{2}}},$
$\mathbf{B}=\frac{2\log(\frac{1}{\delta})[l](b_{n}-a_{n})}{n}+\frac{2\log(\frac{1}{\delta})}{n}\max_{1\leq j\leq m}|R_{nj}|.$
Note that (A.36) implies, for any user-defined $c_{n}\searrow 0$,
@@
$\mathbb{P}\Big(\Big|\frac{1}{n}r_{n}\Big|>c_{n}\Big)\leq 2\exp\left(-2\Big(\frac{(n/[l])c_{n}}{(b_{n}-a_{n})}\Big)^{2}\right).$ (A.39)
(A.37) indicates that, for any user-defined $t_{n}\searrow 0$ and $\theta^{\prime}\in(0,\frac{1}{2})$,
$\mathbb{P}\Big(\max_{1\leq j\leq m}\frac{|R_{nj}|}{\sqrt{n[l]}}>t_{n}\Big)$ $\leq 2m\exp\left(-0.5\Big(\frac{\sqrt{n[l]}t_{n}}{\triangle_{n}\Psi_{n}}\Big)^{2}\right);$ (A.40)
@@
$\mathbb{P}\Big(\max_{1\leq j\leq m}\frac{|R_{nj}|}{n^{\theta^{\prime}}}>\frac{1}{2}\Big)$ $\leq 2m\exp\left(-\Big(\frac{n^{\theta^{\prime}}}{16\triangle_{n}\Psi_{n}}\Big)^{2}\right).$ (A.41)
Combining (A.38)-(A.41), for user-defined $\theta\in(0,\frac{1}{6}]$, $\theta^{\prime}\in(0,\frac{1}{2})$ and $t_{n}\searrow 0$, we have
$\mathbb{P}\Big(|\frac{1}{n}S_{nn}(h_{n})|\leq\tilde{\nu}_{n}(\delta)\max\{\mathbf{A}^{\prime},\mathbf{B}^{\prime}\}+\frac{1}{n^{\frac{1}{2}+\theta}}\Big)\geq 1-2\delta-\sum_{k=1}^{3}\text{Error}_{k},$ (A.42)
$\text{where}\ \mathbf{A}^{\prime}$ $=\sqrt{\frac{2\log(\frac{1}{\delta})[\sum_{j=1}^{m}(\sum_{i\in B_{j}}h_{n}(Z_{ni})-\bar{H}_{n})^{2}]}{n^{2}}}+t_{n}\sqrt{2\log(\frac{1}{\delta})}+\sqrt{\frac{2\xi_{n}\log(\frac{1}{\delta})}{n^{2}}},$
$\mathbf{B}^{\prime}$ $=\frac{2\log(\frac{1}{\delta})[l](b_{n}-a_{n})}{n}+\frac{\log(\frac{1}{\delta})}{n^{1-\theta^{\prime}}},$
and $\text{Error}_{k}$, $k=1,2,3$, coincides with the upper bounds in (A.39)-(A.41).

### v3

@@ A.2 Proof of Results in Section 2
Step 2 According to the proof of Theorems 2 and 3, the following basic triangle inequality holds almost surely.
$\frac{1}{n}|S_{nn}(h_{n})|\leq\frac{1}{n}|S_{nb_{m}}(h_{n})|+\frac{1}{n}|r_{n}|=\frac{1}{n}|\sum_{j=1}^{m}M_{nj}|+\frac{1}{n}|r_{n}|.$
For any user-defined positive $c_{n}$, $t_{n}$ and $s_{n}$, define events
$E_{1}=\Big\{\frac{1}{n}|r_{n}|\leq c_{n}\Big\},\ E_{2}=\Big\{\max_{1\leq j\leq m}\frac{|R_{nj}|}{\sqrt{n[l]}}\leq t_{n}\Big\},\ E_{3}=\Big\{\max_{1\leq j\leq m}\frac{|R_{nj}|}{n}\leq s_{n}\Big\}.$
Then, (A.39) indicates
[…]
$=$ $:\sum_{k=1}^{3}\text{Error}_{k}$ (A.40)
Since $\frac{1}{n}|S_{nn}(h_{n})|\leq\frac{1}{n}|\sum_{j=1}^{m}M_{nj}|+\frac{1}{n}|r_{n}|$ holds almost surely, we thus focus on finding some $U_{n}(\delta)\geq 0$ such that
$\mathbb{P}(\frac{1}{n}|\sum_{j=1}^{m}M_{nj}|+\frac{1}{n}|r_{n}|\leq U_{n}(\delta)|E_{1}\cap E_{2}\cap E_{3})\geq 1-3\delta,\ \ \forall\ \delta>0.$
Please note that, on condition of $E_{1}\cap E_{2}\cap E_{3}$, the following three points hold.
  • B1
    $|\frac{1}{n}r_{n}|\leq c_{n}$;
  • B2
    $\frac{1}{n}|M_{nj}|\leq\frac{1}{n}|T_{nj}|+\frac{1}{n}|R_{nj}|\leq\frac{l}{n}(b_{n}-a_{n})+\frac{1}{n}\max_{1\leq j\leq m}|R_{nj}|\leq\frac{l}{n}(b_{n}-a_{n})+s_{n}$;
  • B3
    According (A.34) and the arbitrary $\xi_{n}$ there, by letting $\xi_{n}=t^{2}_{n}$, we have
    $\sqrt{\frac{\sum_{j=1}^{m}M_{nj}^{2}}{n^{2}}}$ $\leq\sqrt{\frac{\sum_{j=1}^{m}T_{nj}^{2}}{n^{2}}}+\sqrt{\frac{(\max_{1\leq j\leq m}|R_{nj}|)^{2}}{n[l]}}+\frac{t_{n}}{n}\leq\sqrt{\frac{\sum_{j=1}^{m}T_{nj}^{2}}{n^{2}}}+(1+\frac{1}{n})t_{n}$
    $\leq\sqrt{\frac{\sum_{j=1}^{m}(\sum_{i\in B_{i}}(h_{n}(Z_{ni})-\bar{H}_{n}))^{2}}{n^{2}}}+\Big|\frac{1}{n}\sum_{j=1}^{m}M_{nj}\Big|\frac{1}{\sqrt{m}}+(1+\frac{1}{n})t_{n},$
    where the last inequality is due to (A.36).
Thus, based on B1-B3 above, using (9) in Theorem 1 yields
$\mathbb{P}$ $\left(\frac{1}{n}|\sum_{j=1}^{m}M_{nj}|+\frac{1}{n}|r_{n}|\leq U_{n}(\delta)|E_{1}\cap E_{2}\cap E_{3}\right)\geq 1-3\delta,$
$U_{n}(\delta)=$ $\sqrt{\frac{2\log(\frac{1}{\delta})\sum_{j=1}^{m}(\sum_{i\in B_{i}}(h_{n}(Z_{ni})-\bar{H}_{n}))^{2}}{n^{2}}}+\Big|\frac{1}{m}\sum_{j=1}^{m}M_{nj}\Big|\sqrt{\frac{2\log(\frac{1}{\delta})}{m}}$
$+(1+\frac{1}{n})t_{n}\sqrt{2\log(\frac{1}{\delta})}+3.15\log(\frac{1}{\delta})\Big(\frac{l(b_{n}-a_{n})}{n}+s_{n}\Big)+c_{n}.$
Some simple algebra shows
@@
$\mathbb{P}$ $\left(|\frac{1}{n}\sum_{j=1}^{m}M_{nj}|+\tilde{\nu}_{n}(\delta)\frac{|r_{n}|}{n}\leq\tilde{\nu}_{n}(\delta)U^{\prime}_{n}(\delta)\Big|E_{1}\cap E_{2}\cap E_{3}\right)\geq 1-3\delta,\ \ \forall\ \delta>0,$
$U_{n}^{\prime}(\delta)$ $=\sqrt{\frac{2\log(\frac{1}{\delta})\hat{V}_{n}}{n}}++(1+\frac{1}{n})t_{n}\sqrt{2\log(\frac{1}{\delta})}+3.15\log(\frac{1}{\delta})\Big(\frac{l(b_{n}-a_{n})}{n}+s_{n}\Big)+c_{n},$
$\hat{V}_{n}$ $=\frac{1}{n}\sum_{j=1}^{m}(\sum_{i\in B_{i}}(h_{n}(Z_{ni})-\bar{H}_{n}))^{2},\tilde{\nu}_{n}(\delta)=\Big(1-\sqrt{\frac{2\log(1/\delta)}{[n/[l]]}}\Big)^{-1}.$
Additionally, since $\tilde{\nu}_{n}(\delta)>1$ holds strictly, we have
@@
$\mathbb{P}$ $\left(|\frac{1}{n}S_{nn}(h_{n})|\leq\tilde{\nu}_{n}(\delta)U^{\prime}_{n}(\delta)\Big|E_{1}\cap E_{2}\cap E_{3}\right)=:\mathbb{P}(E_{0}|E_{1}\cap E_{2}\cap E_{3})\geq 1-3\delta,\ \ \forall\ \delta>0.$
Finally, using conditional probability formula yields
$\mathbb{P}(E_{0})$ $=\mathbb{P}(E_{1}\cap E_{2}\cap E_{3})\mathbb{P}(E_{0}|E_{1}\cap E_{2}\cap E_{3})$
$\geq(1-\sum_{k=1}^{3}\mathbb{P}(E_{k}^{c}))(1-\delta)\geq(1-\sum_{k=1}^{3}\text{Error}_{k})(1-\delta)$
$\geq 1-\delta-\sum_{k=1}^{3}\text{Error}_{k}.$
We thus finish the proof.